OpenStudy (anonymous):

Suppose f(x)=4x-2 and g(x)=-2x+1. find the value of f(5)/g(-3)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f(x) = 4x - 2 f(5) = 4*5 - 2 f(5) = ???

jimthompson5910 (jim_thompson5910):

g(x)=-2x+1 g(-3)=-2*(-3)+1 g(-3) = ???